the length of a rectangle is 1 inch less than twice that of its width. If the perimeter measures 49 inches, then find the dimensions of the rectangle.
Answer by fractalier(6550)

L = 2W - 1
P = 2L + 2W = 49
Now substituting, we get
2(2W - 1) + 2W = 49
4W - 2 + 2W = 49
Simplifying we get
6W = 51 and
W = 8.5
Then
L = 2W - 1 = 2(8.5) - 1 = 17 - 1 = 16